solve for x in the following equations

  1. 4x-2 = 3x+4
  2. 9x-3x-8=0
1
Expert's answer
2021-03-16T07:14:09-0400

1) Let's take logarithm of both sides of equation:


"log(4^{x-2})=log(3^{x+4}),""(x-2)\cdot log(4)=(x+4)\cdot log(3),""xlog(4)-2log(4)=xlog(3)+4log(3),""xlog(4)-xlog(3)=4log(3)+2log(4),""x(log(4)-log(3))=4log(3)+2log(4),""x(log(\dfrac{4}{3}))=log(81)+log(16),""x(log(\dfrac{4}{3}))=log(1296),""x=\dfrac{log(1296)}{log(\dfrac{4}{3})}."

Answer:

"x=\dfrac{log(1296)}{log(\dfrac{4}{3})}."

2)

"9^x-3^x-8=0,""3^{2x}-3^x-8=0."

Let "3^x=y", then we get the following quadratic equation:


"y^2-y-8=0."

This quadratic equation has two roots:


"y_1=\dfrac{-b+\sqrt{b^2-4ac}}{2a}=\dfrac{1+\sqrt{33}}{2},""y_2=\dfrac{-b-\sqrt{b^2-4ac}}{2a}=\dfrac{1-\sqrt{33}}{2}."

Let's go back, substitute "y" and solve for "x":


"3^x=\dfrac{1+\sqrt{33}}{2},""log(3^x)=log(\dfrac{1+\sqrt{33}}{2}),""xlog(3)=log(\dfrac{1+\sqrt{33}}{2}),""x=\dfrac{log(\dfrac{1+\sqrt{33}}{2})}{log(3)}."

There is no solution for case when we substitute "y_2=\dfrac{1-\sqrt{33}}{2}."

Answer:

"x=\dfrac{log(\dfrac{1+\sqrt{33}}{2})}{log(3)}."